# Variational Scarring in Open Two-Dimensional Quantum Dots

### v2

@@ Abstract
Quantum chaos – sometimes referred to as quantum chaology [1, 2] – explores the transition from quantum states to classical behavior, a topic of intense debate since the advent of quantum mechanics [3, 4]. While classical mechanics exhibits true chaos, quantum systems display signatures of chaotic behavior in distinct ways [5, 6, 7, 8], such as characteristic spectral statistics, wavefunction scarring, and other non-ergodic phenomena. At the heart of this exploration lies the correspondence principle, which is intimately linked to longstanding questions regarding measurement and decoherence in quantum mechanics [9, 10]. One proposed framework for understanding the emergence of classicality is quantum Darwinism, which posits that decoherence – by suppressing most quantum superpositions – yields environment-induced superselection (einselection). This process favors the survival of robust, pointer states that proliferate their information into the environment, while fragile quantum states are suppressed [11]. Intriguingly, these pointer states often exhibit quantum scarring, thus defying the expected ergodic behavior of quantum-chaotic systems [12].
@@
In general, open quantum dots (QDs) provide a versatile platform for investigating quantum scars, offering valuable insights into classical-quantum correspondence partially due to their mixed classical phase space structure. While a single scarred state is typically thought to be fragile against strong perturbations, such as system opening, the proliferation of scars can, however, result in significant robustness for the entire family of scarred states within a moderate energy window. Consequently, these scars can have a substantial effect on the transmission properties of quantum devices [29, 33, 43, 44]: notable transmission variations over small energy scales have been observed, with the fractal conductance fluctuations in quantum devices often attributed to their chaotic nature.
@@
In this Letter, we investigate the quantum transport of a two-dimensional (2D) device illustrated in Fig. 1 in which the coupling to the reservoirs is mediated by a pair of quantum point contacts. We assume that the QD contains impurities introducing randomly positioned bumps within the confining potential, whereas a more controlled, localized disturbance is generated via a nanotip. In particular, the perturbation gives rise to a recently discovered form of scarring within the closed QD [45, 46, 47, 48, 49, 50, 51]: these scars emerge in a synthesis of the quantum near-degeneracy in the unperturbed system and the localized nature of the perturbation. This unique birth mechanism [48] sets the variational scars apart from the conventional, Heller-type scarring [13, 17, 18], extending the concept of scarring into systems lacking suitable classical POs or when those orbits are too unstable to sustain conventional scarring. In other words, this origin allows scars to emerge and thrive under conditions that would typically inhibit conventional scarring. Furthermore, it highlights the fact that the near-degeneracies of a system can have a profound impact, not only the exact degeneracies associated with the continuous symmetries, as well as underscores the subtle role that disorder can play in a quantum device.
In particular, we here demonstrate that variational scars not only persist but also leave a distinct signature in the quantum transmission of an open QD, a fascinating prospect that has remained largely unexplored. We further support this finding by correlating the observed transmission peaks with a local density of states bearing the imprint of a scar, thus paving the way towards a very first experimental observation and successive utilization of variational scarring in a QD device.
@@
$H=\frac{1}{2}\nabla^{2}+V_{\text{ext}}+V_{\text{tip}}+V_{\text{imp}},$ (1)
@@
In addition to the confining potential $V_{\textrm{ext}}$, recognizing that real quantum devices are often affected by disorder arising from impurities and imperfections – such as atoms or ions migrating into a QD system – we model the disorder potential as a sum of randomly distributed Gaussian bumps $V_{\text{imp}}=A_{n}\sum_{i}\exp(-\frac{|\bm{r}-\bm{r}_{i}|^{2}}{2\sigma_{n}^{2}})$, where $A_{n}$ and $\sigma_{n}$ represent the amplitude and width of the impurity bumps, respectively. This kind of disorder mode has been validated for a QD by density-functional theory [68, 69] and the diffusive quantum Monte-Carlo approach [62]. Moreover, we introduce a controllable localized perturbation stemming from a conducting nanotip for which a well-established approximation [70, 71, 72] is also the Gaussian profile $V_{\text{tip}}=A_{T}\exp(-{\frac{|\mathbf{r}-\mathbf{r}_{0}|^{2}}{2\sigma_{T}^{2}}})$, centered at the location $\bm{r}_{0}$, where $A_{T}$ and $\sigma_{T}$ denote its amplitude and width.
@@
As a reference, we first consider the case where there is no nanotip (i.e., $V_{tip}=0$), and the only perturbation comes from the impurity bumps $V_{\textrm{imp}}$, with amplitude $A_{n}=2$ and width $\sigma_{n}\simeq 0.05$. In this scenario, we find that several states accommodate a BB scar. Within this impurity strength regime, the perturbation $V_{\textrm{imp}}$ mostly leads to linear combinations of a single resonant set. Therefore, we can approximate the perturbed eigenstates employing the degenerate perturbation theory, i.e., by diagonalizing $V_{\textrm{imp}}$ within the near-degenerate subspace of the given resonant set. According to the variational principle, these formed states are associated with the extrema of the expectation value of the perturbation. Due to the spatially localized nature of the perturbation, the scarred states are thus favored over non-scarred states whose probability density commonly covers a much larger region of space (see, e.g., Refs. [45, 48])
@@
By the previous argument the scar orientations are mostly selected by the positions of the impurities resulting the most effective externalization of the perturbation: The BB scars shown in Fig. 2(a) orient themselves to maximize or minimize the perturbation by coinciding with or avoiding as many bumps as possible, respectively. While the scars maintain a similar alignment across a broad energy range, controlling the scar direction is challenging due to the inherent randomness of the impurity perturbation. However, by introducing an additional, stronger bump (i.e., $V_{tip}\neq 0$) via the nanotip ($A_{T}=12$ and $\sigma_{T}=0.1$), the BB scars align with the position of the nanotip, as seen in Fig. 2(b), which provides a feasible strategy to design the orientation of the scars experimentally. Furthermore, it is also worth noting that the single perturbation caused by the tip can be sufficient to generate distinct scars [46].
Figure 2: (a) Examples of eigenstates in a closed, perturbed quantum dot containing random impurities (characterized by amplitude $A_{n}=2$ and width $\sigma_{n}\simeq 0.05$). The positions of impurities are represented by green dots. (b) Examples of eigenstates with the same system and quantum numbers $n$ as in (a), but now including a nanotip generating an additional, more prominent potential bump (amplitude $A_{T}=12$, width $\sigma_{T}=0.1$), which is indicated as a red dot.
@@
Nevertheless, the influence of these scars on the electrical transmission of an open quantum device has remained elusive. To address this question, we couple the QD to two leads, effectively opening the system. The leads are modeled as semi-infinite waveguides, with harmonic confinement in the transverse direction and free propagation along the longitudinal axis. This setup ensures smooth continuity and consistent boundary conditions at the interface with the scattering region. The coupling between the leads and the QD is characterized by the rate operator $\Gamma$, which reflects the level broadening induced by the reservoirs. To quantify the coupling strength, we define a dimensionless parameter $\mathcal{C}=\textrm{Tr}[\hat{\Gamma}(\omega)]/\delta$, where $\delta$ is the mean level spacing of the isolated QD in the energy window of interest. This ratio distinguishes between weak coupling ($\mathcal{C}\ll 1$), where transmission is dominated by sharp, isolated resonances, and intermediate to strong coupling ($\mathcal{C}\gtrsim 1$), where significant level broadening leads to overlapping resonances and more delocalized transport pathways.
The transmission through the QD is then determined within the Landauer-Büttiker formalism [74, 75, 76] as
$\mathcal{T}(\omega)=\textrm{Tr}\Big{[}\hat{G}^{R}(\omega)\,\hat{\Gamma}(\omega)\,\hat{G}^{A}(\omega)\,\hat{\Gamma}(\omega)\Big{]},$ (2)
which is computed by employing the TINIE software [77]. Moreover, we compute the local density of states (LDOS) as
$\rho(\bm{r},\omega)=-\textrm{Im}\big{[}\hat{G}^{R}(\bm{r},\omega)\big{]},$ (3)
revealing a map of the spatial distribution of quantum states contributing to the transport at specific probing energies $\omega$.

### v3

@@ Abstract
I Introduction
Quantum chaos – sometimes referred to as quantum chaology [1, 2] – explores the transition from quantum states to classical behavior, a topic of intense debate since the advent of quantum mechanics [3, 4]. While classical mechanics exhibits true chaos, quantum systems display signatures of chaotic behavior in distinct ways (see, e.g., Ref. [5] for a general introduction), such as characteristic spectral statistics [6], wavefunction scarring [7], and other non-ergodic phenomena [8]. At the heart of this exploration lies the correspondence principle, which is intimately linked to longstanding questions regarding measurement and decoherence in quantum mechanics [9, 10]. One proposed framework for understanding the emergence of classicality is quantum Darwinism [11, 12, 13], which posits that decoherence – by suppressing most quantum superpositions – yields environment-induced superselection (einselection). This process favors the survival of robust, pointer states that proliferate their information into the environment, while fragile quantum states are suppressed [13]. Intriguingly, these pointer states often exhibit quantum scarring, thus defying the expected ergodic behavior of quantum-chaotic systems [14].
@@ I Introduction
In general, open quantum dots (QDs) provide a versatile platform for investigating quantum scars, offering valuable insights into classical-quantum correspondence partially due to their mixed classical phase space structure. While a single scarred state is typically thought to be fragile against strong perturbations, such as system opening, the proliferation of scars can, however, result in significant robustness for the entire family of scarred states within a moderate energy window. Consequently, these scars can have a substantial effect on the transmission properties of quantum devices. For instance, non-exponential decay and tunneling anomalies attributed to scarring effects have been experimentally observed in open systems [31, 35]. Similarly, theoretical studies have shown that scarred states can dominate transport in open QDs under certain conditions [45]. Moreover, experimental works on mesoscopic devices have demonstrated strong energy-dependent variations in conductance, suggesting the presence of scar-mediated pathways [46]. Notable transmission variations over small energy scales have also been attributed to the underlying chaotic dynamics of the quantum dot, which give rise to fractal conductance fluctuations [47, 48].
@@
In this article, we investigate the quantum transport of a two-dimensional (2D) device illustrated in Fig. 1 in which the coupling to the reservoirs is mediated by a pair of quantum point contacts. We assume that the QD contains impurities introducing randomly positioned bumps within the confining potential, whereas a more controlled, localized disturbance is generated via a nanotip. In particular, the perturbation gives rise to a recently discovered form of scarring within the closed QD [49, 50, 51, 52, 53, 54, 55]: these scars emerge in a synthesis of the quantum near-degeneracy in the unperturbed system and the localized nature of the perturbation. This unique birth mechanism [52] sets the variational scars apart from the conventional, Heller-type scarring [15, 19, 20], extending the concept of scarring into systems lacking suitable classical POs or when those orbits are too unstable to sustain conventional scarring. In other words, this origin allows scars to emerge and thrive under conditions that would typically inhibit conventional scarring. Furthermore, it highlights the fact that the near-degeneracies of a system can have a profound impact, not only the exact degeneracies associated with the continuous symmetries, as well as underscores the subtle role that disorder can play in a quantum device.
In particular, we demonstrate here that variational scars not only persist but also leave a distinct signature in the quantum transmission of an open QD, a fascinating prospect that has remained largely unexplored. We further support this finding by correlating the observed transmission peaks with a local density of states (LDOS) [56] bearing the imprint of a scar, thus paving the way towards a very first experimental observation and successive utilization of variational scarring in a QD device.
II System
@@
$\bm{H}_{QD}=\frac{1}{2}\nabla^{2}+V_{\text{ext}}+V_{\text{tip}}+V_{\text{imp}},$ (1)
@@
In addition to the confining potential $V_{\textrm{ext}}$, recognizing that real quantum devices are often affected by disorder arising from impurities and imperfections – such as atoms or ions migrating into a QD system – we model the disorder potential as a sum of randomly distributed Gaussian bumps
$V_{\text{imp}}=A_{n}\sum_{i}\exp\left(-\frac{|\bm{r}-\bm{r}_{i}|^{2}}{2\sigma_{n}^{2}}\right),$ (2)
where $A_{n}$ and $\sigma_{n}$ represent the amplitude and width of the impurity bumps, respectively. This kind of disorder mode has been validated for a QD by density-functional theory [73, 74] and the diffusive quantum Monte-Carlo approach [67]. Moreover, we introduce a controllable localized perturbation stemming from a conducting nanotip for which a well-established approximation [75, 76, 77] is also the Gaussian profile
$V_{\text{tip}}=A_{T}\exp\left(-{\frac{|\mathbf{r}-\mathbf{r}_{0}|^{2}}{2\sigma_{T}^{2}}}\right),$ (3)
centered at the location $\bm{r}_{0}$, where $A_{T}$ and $\sigma_{T}$ denote its amplitude and width.
III Variational Scarring in Closed Quantum Dot
Figure 2: (a) Selected eigenstates of a closed, perturbed quantum dot containing both random Gaussian impurities (green dots) with average amplitude $A_{n}=2$ and width $\sigma_{n}\simeq 0.05$, as well as a nanotip modeled as a Gaussian bump (red dot) with amplitude $A_{T}=12$ and width $\sigma_{T}=0.1$. These eigenstates include the first three clearly identifiable bouncing-ball scars (the middle states) and adjacent states in energy. (b) Corresponding eigenstates (same quantum numbers $n$) for the system with only random Gaussian impurities, illustrating the absence of prominent directional localization, and thereby highlighting the role of the nanotip in inducing the scarred patterns. The quantum numbers and corresponding eigenenergies are indicated in the figure.
@@ III Variational Scarring in Closed Quantum Dot
As a reference, we first consider the case where there is no nanotip (i.e., $V_{tip}=0$), and the only perturbation comes from the impurity bumps $V_{\textrm{imp}}$, with amplitude $A_{n}=2$ and width $\sigma_{n}\simeq 0.05$. In this scenario, we find that several states accommodate a BB scar. Within this impurity strength regime, the perturbation $V_{\textrm{imp}}$ mostly leads to linear combinations of a single resonant set. Therefore, we can approximate the perturbed eigenstates employing the degenerate perturbation theory, i.e., by diagonalizing $V_{\textrm{imp}}$ within the near-degenerate subspace of the given resonant set. According to the variational principle, these formed states are associated with the extrema of the expectation value of the perturbation. Due to the spatially localized nature of the perturbation, the scarred states are thus favored over non-scarred states whose probability density commonly covers a much larger region of space (see, e.g., Refs. [49, 52]).
@@
By the previous argument the scar orientations are mostly selected by the positions of the impurities resulting the most effective externalization of the perturbation: The BB scars shown in Fig. 2(b) orient themselves to maximize or minimize the perturbation by coinciding with or avoiding as many bumps as possible, respectively. While the scars maintain a similar alignment across a broad energy range, controlling the scar direction is challenging due to the inherent randomness of the impurity perturbation. However, by introducing an additional, stronger bump (i.e., $V_{tip}\neq 0$) via the nanotip ($A_{T}=12$ and $\sigma_{T}=0.1$), the BB scars align with the position of the nanotip, as seen in Fig. 2(a), which provides a feasible strategy to design the orientation of the scars experimentally.
Intriguingly, variational scarring appears to persist even up to the semiclassical limit $\hbar\rightarrow 0$, owing to the survival of resonant sets, unless destroyed by decoherence. Analogously, non-ergodic BB states in stadium billiards are proven to linger in this limit [79]. However, it remains as an open problem whether variational scars form a zero-measure subset in the semiclassical limit, as is the case for their Heller-type cousins, ergo aligning with the ergodicity theorems of Refs. [18, 16, 17]. Moreover, this onset of ergodicity can be exceedingly slow, as exemplified by the study of a particle in a tilted box [80].
At the same time, the total perturbation responsible for scarring must be strong enough to mix eigenstates within a resonant set, yet weak enough to avoid coupling these states to the broader spectrum, which would otherwise result in chaotic-looking combinations [51]. As $\hbar\rightarrow 0$, the energy levels may become denser, and the resonant set may thus lose its isolation. Consequently, the same perturbation strength begins to mix a wide range of unperturbed states, thereby suppressing scarring in the high-energy regime.
Nonetheless, variational scars are not rare phenomena; In fact, they occur thrive across various disordered nanostructures. For instance, scars have been demonstrated in a broad range of power-law confinement potentials $V_{\textrm{ext}}\propto r^{a}$ ($a\in\mathbb{R}$[49, 51], as well as in a Rydberg atom set-up [53]. While homogeneousness of a confinement simplify the classical POs structure, variational scars emerge in many other types of potentials landscapes, such as non-homogeneous hyperbolic cosine wells [49, 51] or anisotropic oscillators [52]. A necessary element is the presence of a resonant set associated with classical POs that local perturbations can effectively mix up to form scars. Notable, exact symmetry is not essential; near-symmetries are sufficient. Indeed, scars have been demonstrated [51, 52] in power-law potentials with non-integer exponents $\alpha$, and in elliptical oscillators or power potentials subjected to perpendicular magnetic fields. In all aforementioned cases, closed classical orbits are absent.
On the other hand, the variational mechanism [52] is agnostic to whether the perturbative bumps are attractive or repulsive. While our discussion focus on repulsive perturbations, scars also arising from attractive bumps, i.e., potential dips, that has also been confirmed by prior studies [52, 55]. Generally, lower bump density allows the eigenstates to reflect more of the symmetry of unperturbed system, whereas higher density leads to delocalization and the disappearance of scarring. Similar effects arise from variations in bump amplitude and width. The specific behavior depends on multiple factors [51, 52]: bump parameters (density, amplitude, width), the degree of near-degeneracy in the unperturbed spectrum, and external influences such as applied magnetic fields, hence providing several avenues for controlling scarring.
Furthermore, even in the absence of impurities, a single localized perturbation, such as from a nanotip, can be sufficient to induce distinct scars [50]. However, as shown in Ref. [50], the exact shape of the scar is then highly sensitive to the tip location, while the orientation and geometry of scars caused by multiple perturbations remain robust [49, 51]. For example, pinned pentagram scars adjust their self-crossing point in response to radial movement of the tip, whereas BB scars, which are the primary center of attention here, exhibit robustness but can be reoriented with the tip, as also illustrated in Ref. [55] (see also Fig. 2).
IV Quantum Transport Methodology
@@
Nevertheless, the influence of these scars on the electrical transmission of an open quantum device has remained elusive. To address this question, we couple the QD to two leads, effectively opening the system. All transport and related quantities are numerically computed utilizing the TINIE software package [81], which we briefly outline below.
The leads are modeled as semi-infinite waveguides, with harmonic confinement in the transverse direction and free propagation along the longitudinal axis. This setup ensures smooth continuity and consistent boundary conditions at the interface with the scattering region. Within the Landauer–Büttiker framework [82, 56, 83], the influence of the given lead $\alpha$ on the central region is encapsulated in the retarded and advanced self-energies, i.e., $\bm{\Sigma}_{\alpha}^{R}(\omega)$ and $\bm{\Sigma}_{\alpha}^{A}(\omega)$, respectively:
$\bm{\Sigma}^{R}(\omega)=$ $\bm{V}^{\dagger}_{\alpha}\;[(\omega+i\eta\mathbbm{1})-\bm{H}_{L_{\alpha}}]^{-1}\;\bm{V}_{\alpha}\;,$ (4)
$\bm{\Sigma}^{A}(\omega)=$ $\bm{V}^{\dagger}_{\alpha}\;[(\omega-i\eta\mathbbm{1})-\bm{H}_{L_{\alpha}}]^{-1}\;\bm{V}_{\alpha}\;,$ (5)
where $\bm{H}_{L_{\alpha}}$ is the isolated lead $\alpha$ Hamiltonian, $\mathbbm{1}$ is the identity matrix in the lead subspace, and $\eta>0$ is an infinitesimal parameter that fixes the boundary conditions.
The matrix $\bm{V}_{\alpha}$ describes the coupling between the lead–dot interface slice of lead $\alpha$ and the adjacent sites in QD. It is constructed by projecting the full device Hamiltonian $\bm{H}_{\text{tot}}$ onto the basis functions of the lead and the dot. In block form:
$\bm{H}_{\text{tot}}=\begin{pmatrix}\bm{H}_{L_{1}}&\bm{V}_{L_{1}}&0\\ \bm{V}_{L_{1}}^{\dagger}&\bm{H}_{\text{QD}}&\bm{V}_{L_{2}}^{\dagger}\\ 0&\bm{V}_{L_{2}}&\bm{H}_{L_{2}}\end{pmatrix},$ (6)
where $\bm{H}_{\text{QD}}$ is the isolated-dot Hamiltonian, $\bm{H}_{L_{1/2}}$ are the lead Hamiltonians, and $\bm{V}_{1/2}$ are the lead–dot coupling blocks. Accordingly, the elements of $\bm{V}_{\alpha}$ are
$[\bm{V}_{\alpha}]_{ij}=\langle\phi_{\alpha,i}|\bm{H}_{\text{tot}}|\psi_{\text{QD},j}\rangle,$ (7)
with $\phi_{\alpha,i}$ a basis function on the interface slice of the lead $\alpha$, and $\psi_{\text{QD},j}$ a basis function in the dot region.
The coupling strength of the lead–dot contact is quantified by the energy-dependent total rate operator $\bm{\Gamma}_{\text{tot}}(\omega)=\sum_{\alpha}\bm{\Gamma}_{\alpha}(\omega)$ with the lead-specific rates
$\bm{\Gamma}_{\alpha}=i[\bm{\Sigma}_{\alpha}^{R}-\bm{\Sigma}_{\alpha}^{A}],$ (8)
reflecting the level broadening induced by coupling to an individual lead $\alpha$. Since $\textrm{Tr}[\bm{\Gamma}_{\text{tot}}(\omega)]$ represents the total broadening of QD energy levels due to coupling with the leads, we define a dimensionless coupling parameter as
$\mathcal{C}=\textrm{Tr}[\bm{\Gamma}_{\text{tot}}(\omega)]/\delta$ (9)
where $\delta$ is the mean level spacing of the isolated QD within the energy window of interest. This ratio distinguishes between weak coupling ($\mathcal{C}\ll 1$), where transmission is dominated by sharp, isolated resonances, and intermediate to strong coupling ($\mathcal{C}\gtrsim 1$), where significant level broadening leads to overlapping resonances and more delocalized transport pathways.
The transmission between two distinct leads $\alpha$ and $\beta$, through the QD is then determined as
$\mathcal{T}_{\alpha\beta}(\omega)=\textrm{Tr}\Big{[}\bm{G}^{R}(\omega)\,\bm{\Gamma}_{\beta}(\omega)\,\bm{G}^{A}(\omega)\,\bm{\Gamma}_{\alpha}(\omega)\Big{]},$ (10)
where $\hat{G}^{R}(\omega)$ and $\hat{G}^{A}(\omega)$ denote the retarded and advanced Green’s functions of the open quantum system, respectively [82]. These Green’s functions incorporate the effect of coupling to the leads through the retarded self-energies $\Sigma_{\alpha}^{R}(\omega)$, and are defined as
$\bm{G}^{R}(\omega)$ $=\left[\omega\mathbbm{1}-\bm{H}_{\text{QD}}-\sum_{\alpha}\Sigma_{\alpha}^{R}(\omega)\right]^{-1}=[\bm{G}^{A}(\omega)]^{\dagger}.$ (11)
In addition to transmission coefficient of Eq. 10, we compute LDOS to resolve the spatial distribution of quantum states contributing to transport at specific probe energies $\omega$. As defined in the context of Green’s function formalism [83], the LDOS is given by
$\rho(\bm{r},\omega)=-\textrm{Im}\big{[}\bm{G}^{R}(\bm{r},\omega)\big{]},$ (12)
where $\bm{G}^{R}(\bm{r},\omega)$ is subsequently the projection of the retarded Green’s function in real space [56], evaluated at position $\bm{r}$ as
$\bm{G}^{R}(\bm{r},\omega)=\sum_{i,j}\psi_{\text{QD},i}^{*}(\bm{r})\,[\bm{G}^{R}(\omega)]_{ij}\,\psi_{\text{QD},j}(\bm{r}).$ (13)
This measure serves as a powerful diagnostic, opening up a way to correlate transport features of the open system with the underlying scarred wavefunction structure of the closed counterpart.
V Transport Signatures of Variational Scars
Figure 3: Logarithm of transmission vs. probe energy within bias window 2324232423–2423 – 24 for noise-only case with $\mathcal{C}\simeq 0.001$ (blue) and nanotip+noise with $\mathcal{C}\simeq 0.002$ (orange), both in weak coupling limit ($\mathcal{C}\ll 1$). The strongest peaks are labeled as (P1) for noise-only and (P2) for nanotip+noise. Logarithmic scale is used due to the wide range of transmission values. Below, normalized LDOS maps are shown at the corresponding peak energies (P1) and (P2) in panels (a) and (b), respectively, illustrating distinct bouncing-ball scar pattern induced by the tip. Panels (c)-(e) display normalized LDOS of nanotip+noise system at the energy corresponding to strongest transmission peak in weakly coupled leads ($\omega=23.350$), now under stronger coupling strengths $\mathcal{C}$: (c) $\simeq 7.55$, (d) $\simeq 48.34$, and (e) $\simeq 121.20$.
Figure 4: Logarithm of transmission vs. probe energy within bias window 2324232423–2423 – 24, shown for intermediate ($\mathcal{C}\simeq 12.21$, blue), and strong ($\mathcal{C}\simeq 96.59$, orange) coupling strengths. Prominent transmissions labeled as (P1) at $\omega=23.127$ and (P2) at energy $\omega=23.629$. Below, the corresponding normalized LDOS maps at these peak energies are shown in panels (a) and (b), respectively, revealing how increased self-energy disrupts the scarred patterns observed at weaker coupling, resulting in more extended wavefunction distributions.